OpenStudy (anonymous):

what is 7/8 divided by 2/3

OpenStudy (anonymous):

reverse the second fraction and multiply \[\frac{ 7 }{ 8 }*\frac{ 3 }{ 2 } = \frac{ 21 }{ 16 }\]

OpenStudy (anonymous):

no not 3/2 its 2/3

OpenStudy (anonymous):

do 0.46 come before 1.9

OpenStudy (anonymous):

Hi hotdog. First...he multiplied 7/8 and 3/2 because when you divide 2 fractions, you multiply by the inverse.( You flip the second fraction over to multiply) .46 is less than 1.9.